OpenStudy (anonymous):

HELP!!!!! Which equation matches the table of values? x y 0 1 1 1/4 2 1/16 3 1/64 y = 4^x y = x^4 y = x y = 4x

OpenStudy (qwertty123):

none of those

OpenStudy (anonymous):

how?

OpenStudy (qwertty123):

y = 4^(-x) = 1/(4^x)

OpenStudy (anonymous):

So should I just pick the 1/4

OpenStudy (anonymous):

oops I meant to type 1/4x

OpenStudy (qwertty123):

Um.. I guess